On June 30, 2021, the Esquire Company sold merchandise to a customer and accepted a noninterest-bearing note in exchange. The note requires payment of $28,000 on March 31, 2022. The fair value of the merchandise exchanged is $26,530. Esquire views the financing component of this contract as significant. 


Required:

 1. Prepare journal entries to record the sale of merchandise (omit any entry that might be required for the cost of the goods sold), any December 31, 2021 interest accrual, and the March 31, 2022 collection.

 2. What is the effective interest rate on the note?
Answer #1
Answer

1)

Date Account title Debit credit
June 30 2021 Note receivable $     28,000
Discount on note receivable $       1,470
sales revenue $     26,530
December 31 2021 Discount on note receivable $          980
Interest revenue (1470*6/9 ) $          980
march 31 2022 Discount on note receivable $          490
Interest revenue (1470-980) $          490
March 31 2022 cash $     28,000
Note receivable $     28,000
2)Discount amount = Face value of note *r *n/12
   1470 = 28000* r *9/12
   1470/28000 = r*9/12
    .0525*12/9 = r
    r = .07 or 7%
Answer #2

The answer has been presented in the supporting sheet. For detailed answer refer to the supporting sheet.

Debit Credit 265301 Answer Part1) Journal Entry Date Account Title June 30,2021. Note Receivable Sales (To record sale of goo